Question: Consider a bag containing 5 green, 7 black and 4 white marbles. If three marbles are drawn randomly then what is the probability that none of them is white? О a. 1021 O b. 11/28 O c. 12/23 o d. 13/18 e.
Answer #1

Total number of marbles = 5 + 7 + 4 = 16

Number of non white marbles = 5+7 = 12

P(none of the 3 marbles drawn are white) = Number of ways to draw 3 non white marbles from 12 / Number of ways to draw 3 marbles from total 16

= 12C3 / 16C3

= 220/560

= 11/28

OR

P(none of the 3 marbles drawn are white) = P(first marbles is not white) x P(second marble from remaining 15 is also not white) x P(third marble from remaining 14 is non white)

= 12/16 x 11/15 x 10/14

= 11/28